# Reconstructing jet anisotropies with cumulants

### v1

@@ Abstract
In relativistic heavy-ion collisions, where quark-gluon plasma (QGP) forms, hadron production is anisotropic at both low and high transverse momentum, driven by flow dynamics and spatial anisotropies. To better understand these mechanisms, we use multi-particle correlations to reconstruct jet anisotropies. We simulate data using TennGen as a hydro-like background and combine it with PYTHIA-generated jets, clustering them with the anti-$k_{T}$ algorithm. Jet anisotropies are unfolded using a Bayesian technique, ensuring the robustness of the reconstructed signals. Our results demonstrate that multi-particle cumulant methods can accurately capture the differential jet azimuthal anisotropies, providing crucial insights into high-$p_{T}$ behavior and the dynamics within heavy-ion collisions.
@@ I Introduction
Relativistic heavy-ion collisions create an extremely hot and dense phase of matter known as the quark-gluon plasma (QGP). Investigations at the Relativistic Heavy Ion Collider (RHIC) Adams et al. (2005); Adcox et al. (2005); Arsene et al. (2005); Back et al. (2005) and the Large Hadron Collider (LHC) Acharya et al. (2024); Hayrapetyan et al. (2024); Busza et al. (2018) have demonstrated a wide array of complex phenomena related to the QGP. A key signature of QGP formation is the azimuthal anisotropy observed in the production of hadrons, particularly at low transverse momentum ($p_{T}$ $<$ 3 GeV). This anisotropy can be effectively modeled as hydrodynamic flow, with the final momentum distribution of hadrons reflecting initial spatial anisotropies, which are translated into momentum space through pressure gradients Hirano et al. (2006); Huovinen et al. (2001); Hirano and Tsuda (2002); Romatschke and Romatschke (2007); Song et al. (2011); Schenke et al. (2011); Lacey et al. (2016). These azimuthal anisotropies are described using a Fourier expansion Voloshin and Zhang (1996); Poskanzer and Voloshin (1998):
$\frac{dN}{d\phi}$ $\propto$ $\left[1+2\sum_{n=1}^{\infty}v_{n}\cos(n(\phi-\psi_{n}))\right]$ (1)
@@
At high transverse momentum ($p_{T}$ $>$ 10 GeV), hadron production in ion-ion (AA) collisions shows significant suppression compared to expectations based on proton-proton ($p$+$p$) collisions, a phenomenon understood as jet quenching Connors et al. (2018); Qin and Wang (2015). In this process, high-energy partons lose energy through both radiative and collisional interactions within the QGP Qin and Wang (2015); Mehtar-Tani et al. (2013); Blaizot and Mehtar-Tani (2015). These high-$p_{T}$ hadrons and their associated jets also exhibit non-zero azimuthal anisotropy, even though they fall outside the typical regime where the hydrodynamic flow would apply Sirunyan et al. (2018); Aad et al. (2013a); Acharya et al. (2018). Instead, these anisotropies are understood to arise from the spatial inhomogeneities in the QGP, with jet quenching effects being more pronounced for partons traveling through longer paths in the QGP Gyulassy et al. (2001); Shuryak (2002); Jia (2013). Both low- and high-$p_{T}$ hadrons, therefore, share a standard orientation of their anisotropies, correlating with the geometry of the colliding nucleons. A long-standing theoretical challenge has been to simultaneously explain both high-$p_{T}$ suppression and the corresponding azimuthal anisotropy observed in A+A collisions and several models have been proposed to address this issue Shuryak (2002); Molnar and Sun (2013); Noronha-Hostler et al. (2016); Betz et al. (2017); Holtermann et al. (2024); Barreto et al. (2022); Zhang and Liao (2013).
@@
Jet production in the QGP is also azimuthally anisotropic, where the anisotropy in Eq. 1 is due to the pathlength dependence of jet quenching. The pathlength is minimized when $\phi$ and $\psi_{n}$ are approximately aligned, leading to less jet attenuation. In contrast, when $\phi$ and $\psi_{n}$ are approximately orthogonal, the pathlength is maximized, which leads to increased jet attenuation. Previous measurements of differential azimuthal anisotropies $v^{\prime}_{n}$ have used an event plane method where the event plane is defined by soft particles  Aad et al. (2013a); Adam et al. (2016a). Given observations in the soft sector, these jet anisotropy coefficients should also be sensitive to the measurement technique. Some studies indicate that the difference could be up to 10% Andres et al. (2020); Zigic et al. (2022a, b), which is on the order of the uncertainty on measurements by ATLAS Aad et al. (2022).
@@
Measurements from smaller collision systems, such as $p$+$p$ and $p$+Pb at the LHC Aad et al. (2013b); Khachatryan et al. (2010, 2015); Abelev et al. (2013a), and $p$+Au, $d$+Au, and 3He+Au at RHIC Aidala et al. (2019); Abdulhamid et al. (2023), have also revealed significant azimuthal anisotropies at low-$p_{T}$, with patterns similar to those observed in heavy-ion collisions. These findings suggest that even smaller systems may produce short-lived droplets of the QGP; indeed, hydrodynamic models have successfully described the low-$p_{T}$ behavior in such systems. However, when it comes to high-$p_{T}$ behavior in smaller collision systems, the expected signatures of jet quenching have not been observed. While some measurements indicate that no jet suppression at high transverse momentum ($p_{T}$) in $p$+Pb and $d$+Au collisions Adare et al. (2016a); Acharya et al. (2020), other studies indicate there may be some suppression Abdulameer et al. (2025). No alterations in the momentum balance of dijets or hadron-jet pairs have been detected. The ATLAS experiment has also reported non-zero azimuthal anisotropies for hadrons with $p_{T}$ up to 12 GeV, suggesting an extension of anisotropy effects into the high-$p_{T}$ regime typically associated with jet quenching. It is unclear if differential jet quenching contributes to the observed high-$p_{T}$ anisotropies Aad et al. (2023). This leaves two related unresolved puzzles: the lack of jet quenching observed in p+A collisions and the mechanism that produces high-$p_{T}$ hadron anisotropies in the absence of jet quenching.
@@
Further investigation into the mechanism responsible for high-$p_{T}$ hadron anisotropies requires two key approaches: (i) comparing data models with both theory and experiment using consistent techniques and (ii) expanding measurements to incorporate multi-particle correlations. Discrepancies between event plane and cumulant methods can reach up to 10% Andres et al. (2020); Zigic et al. (2022a, b); Abdallah et al. (2022). These variations can complicate the interpretation of experimental measurements made with the event plane method. Several factors, such as long-range non-flow effects and initial/final state fluctuations, can influence the flow correlations measured with the two-particle correlation method, potentially amplifying the observed flow signal. Therefore, applying two- and multi-particle cumulant methods in data analysis and theoretical calculations is crucial for a deeper understanding of the mechanisms behind high-$p_{T}$ hadron anisotropies.
@@
In this work, we demonstrate the feasibility of measuring the jet azimuthal anisotropies using the two- and four-particle cumulant method using TennGen Hughes et al. (2022); Mengel et al. (2023); Hughes and Mengel (2019) and PYTHIA Sjostrand et al. (2006, 2008).
@@ II.1 Models
Events are simulated in TennGen Hughes et al. (2022); Hughes and Mengel (2019); Mengel and Hughes (2022), a data-driven background generator designed to reproduce correlations arising from flow. In this model, the $n^{\text{th}}$-order event plane is set to zero for even values of $n$, while for odd values of $n$, the event plane is chosen randomly. TennGen multiplicities and charged particle ratios match experimentally measured values Aamodt et al. (2010); Abelev et al. (2013b). Similar constraints are also applied to the transverse momentum ($p_{T}$) spectra of charged pions, kaons, and protons through random sampling from an initial distribution fitted to a Blast Wave Ristea et al. (2013); Schnedermann et al. (1993) distribution. In TennGen, the generated particle $v^{\prime}_{n}$($\eta$) is sampled from a flat distribution, an acceptable approximation for the region $|\eta|<1.1$. The $v^{\prime}_{n}(p_{T})$ are modeled using a polynomial fit to the measured $v^{\prime}_{n}(p_{T})$ Adare et al. (2016b); Adam et al. (2016b). TennGen is a hydro-like background for jet interactions. The jet signal is generated using PYTHIA 8.307 Sjostrand et al. (2008) with the Monash Tune 13 Skands et al. (2014) and a minimum hard scattering transverse momentum cutoff of $p_{T,\text{hard}}^{\text{min}}>5$ GeV. The leading PYTHIA jets are realigned to simulate a $\frac{dN_{\text{jet}}}{d\phi_{\text{jet}}}$ distribution with non-zero differential azimuthal anisotropies $v^{\prime}_{n,\text{jet}}$. Various functional forms are considered for the momentum dependence of $v^{\prime}_{n,\text{jet}}$, referred to as rotation classes. A total of $10$ million TennGen events are generated for each sample.
II.2 Jet reconstruction
@@
The merged PYTHIA jet constituents and TennGen charged particles are clustered into anti-$k_{T}$ jets with jet resolution parameter $R=0.4$ using the energy recombination scheme implemented with FastJet 3.4.0 Cacciari et al. (2012). The charged final-state hadrons from PYTHIA p+p events are clustered separately before merging with TennGen, which defines the truth jet momentum. The p+p event is used if there is at least one jet with $p_{T,jet}>$ 10 GeV. Ghost particles with $p_{T,\text{ghost}}=1$ MeV are used to estimate the area of clustered merged jets as described in the FastJet manual Cacciari et al. (2012).
@@
The momentum of the merged jets is corrected using the jet-multiplicity background subtraction method Mengel et al. (2023)
$p_{T,jet}^{\text{corr, N}}=p_{T,jet}^{\text{raw}}-\rho_{\text{mult}}(N_{\text{tot}}-\langle N_{\text{PYTHIA}}\rangle).$ (2)
@@
where $N_{\text{tot}}$ is the observed number of particles within the jet, $\langle N_{\text{PYTHIA}}\rangle$ is the average number of particles in a PYTHIA jet of a given matched uncorrected jet momentum bin, and $\rho_{\text{Mult}}$ is the mean transverse momentum per background particle. Merged jet candidates satisfying $p_{T,jet}^{\text{corr}}>5$ GeV and $A_{\text{jet}}>0.6\pi R^{2}$ Steffanic et al. (2023) are geometrically matched back to the jet axis of the PYTHIA signal jet. The closest merged jet to is considered a match if $\Delta R<0.75R$ and the corresponding PYTHIA jet momentum $p_{T,jet}^{\text{PYTHIA}}$ is taken to be the truth momentum.
II.3 Cumulant method
@@
The traditional cumulant method Bilandzic et al. (2011); Jia et al. (2017), is used to construct the $v^{\prime}_{n}\{2k\}(p_{T})$, where $k$ is a positive integer. In this work, the $v^{\prime}_{n}\{2k\}(p_{T})$ is given as:
$v^{\prime}_{n,\text{jet}}\{2\}(p_{T,jet})$ $=$ $\dfrac{\langle 2^{\prime}_{n}\rangle}{\left(\langle 2_{n}\rangle\right)^{1/2}},$
$\langle 2^{\prime}_{n}\rangle$ $=$ $\langle\langle e^{\textit{i}n(\varphi^{A}_{1}(p_{T,jet})-\varphi^{C}_{2})}\rangle\rangle,$
$\langle 2_{n}\rangle$ $=$ $\langle\langle e^{\textit{i}n(\varphi^{C}_{1}-\varphi^{C}_{2})}\rangle\rangle$ (3)
$v^{\prime}_{n,\text{jet}}\{4\}(p_{T,jet})$ $=$ $\dfrac{\langle 4^{\prime}_{n}\rangle}{\left(-\langle 4_{n}\rangle\right)^{3/4}},$
$\langle 4^{\prime}_{n}\rangle$ $=$ $\langle\langle e^{\textit{i}n(\varphi^{A}_{1}(p_{T,jet})+\varphi^{C}_{2}-\varphi^{C}_{3}-\varphi^{C}_{4})}\rangle\rangle$
-- $2~{}\langle\langle e^{\textit{i}n(\varphi^{C}_{1}-\varphi^{C}_{2})}\rangle\rangle\langle\langle e^{\textit{i}n(\varphi^{A}_{1}(p_{T,jet})-\varphi^{C}_{2})}\rangle\rangle,$
$\langle 4_{n}\rangle$ $=$ $\langle\langle e^{\textit{i}n(\varphi^{C}_{1}+\varphi^{C}_{2}-\varphi^{C}_{3}-\varphi^{C}_{4})}\rangle\rangle$ (4)
-- $2~{}\langle\langle e^{\textit{i}n(\varphi^{C}_{1}-\varphi^{C}_{2})}\rangle\rangle\langle\langle e^{\textit{i}n(\varphi^{C}_{1}-\varphi^{C}_{2})}\rangle\rangle.$
@@
where $\varphi^{A(C)}$ is the azimuthal angle of particles in the region $A$($C$) and $\langle\langle 2_{n}\rangle\rangle$ is the two particle correlations of order $n$. The integrated $\varphi$ is coming from region $C$ with $5.13.1<|\eta_{C}|<$ 5.1 and the particle of interest (i.e., jet victor) in the region $|\eta_{\text{jet}}|<1.1-R$. More details about the traditional cumulant method can be found in Appendix A.
II.4 Unfolding
@@
The differential azimuthal anisotropies for jets $v^{\prime}_{n,\text{jet}}\{2k\}(p_{T,jet})$ are binned as a function of uncorrected jet momentum. Jet measurements with $p_{T,jet}$ dependencies are subject to bin-by-bin migration from smearing due to finite resolution in the uncorrected measurement. This smearing can skew the event-averaged differential particle correlations if the functional form of $v^{\prime}_{n,\text{jet}}$ is dependent on jet momentum. If the $v^{\prime}_{n,\text{jet}}$ are independent of jet momenta, then this smearing will not change the results. Measurements of $v^{\prime}_{n,\text{jet}}$ indicate there may be a nonzero dependence on $p_{T,jet}$ meaning that measurement using a cumulant method will need to be unfolded Adam et al. (2016a); Aad et al. (2013a, 2022) We unfold the reconstructed differential single-event $n^{\mathrm{th}}$-order $2k$-particle correlations using the Bayesian unfolding method D’Agostini (1995) in RooUnfold 2.0.0 Brenner et al. (2020).
@@
We construct a $4$-dimensional response matrix using PYTHIA jets (truth jets) matched to PYTHIA +TennGen jets (reconstructed jets). The momentum and azimuthal angle of the PYTHIA jet are taken as the truth momentum, $p_{T,jet}^{\text{truth}}\equiv p_{T,jet}^{\text{PYTHIA}}$ and the azimuthal angle $\phi_{\text{jet}}^{\text{PYTHIA}}$ is used to construct the truth single event correlation $\langle(2k)^{\prime}_{n}\rangle^{\text{truth}}\equiv\langle(2k)^{\prime}_{n}\rangle^{\text{PYTHIA}}$. We then unfold our reconstructed jet differential single-event $n^{\mathrm{th}}$-order $2k$-particle correlations, which has has no matching criteria for the $\langle(2k)^{\prime}_{n}\rangle^{\text{reco}}$. The unfolding procedure is repeated until the change in $\chi^{2}$ between the unfolded and truth $\langle(2k)^{\prime}_{n}\rangle$ becomes less than the uncertainties of the measured $\langle(2k)^{\prime}_{n}\rangle$. This typically occurs between 3-5 iterations, and more iterations beyond this threshold would yield diminishing increases in unfolded $\langle(2k)^{\prime}_{n}\rangle$ resolution. The average value of the unfolded single event differential correlation $\langle(2k)^{\prime}_{n}\rangle^{\text{corr}}$ is calculated for each jet momentum bin and used to compute the differential second and fourth-order cumulants.

### v2

# Reconstructing Jet Azimuthal Anisotropies with Cumulants

@@ Abstract
In relativistic heavy-ion collisions, where quark-gluon plasma forms, hadron production is anisotropic at both low and high transverse momentum, driven by flow dynamics and spatial anisotropies. To better understand these mechanisms, we use multi-particle correlations to reconstruct jet anisotropies. We simulate data using TennGen as a hydro-like background and combine it with Pythia-8 generated jets, clustering them with the anti-$k_{\mathrm{t}}$ algorithm. Jet anisotropies are unfolded using a Bayesian technique, ensuring the robustness of the reconstructed signals. Our results demonstrate that multi-particle cumulant methods can accurately capture the differential jet azimuthal anisotropies, providing crucial insights into high-$p_{T}$ behavior and the dynamics within heavy-ion collisions.
@@ I Introduction
Relativistic heavy-ion collisions create an extremely hot and dense phase of matter known as the quark-gluon plasma (QGP). Investigations at the Relativistic Heavy Ion Collider (RHIC) [19, 22, 25, 26] and the Large Hadron Collider (LHC) [15, 39, 33] have demonstrated a wide array of complex phenomena related to the QGP. A key signature of QGP formation is the azimuthal anisotropy observed in the production of hadrons, particularly at low transverse momentum ($p_{T}$ << 3 GeV). This anisotropy can be effectively modeled as hydrodynamic flow, with the final momentum distribution of hadrons reflecting initial spatial anisotropies, which are translated into momentum space through pressure gradients [40, 45, 41, 63, 71, 64, 50]. These azimuthal anisotropies are described using a Fourier expansion [73, 59]:
$\frac{dN}{d\phi}$ $\propto$ $C\left[1+2\sum_{n=1}^{\infty}v_{n}\cos(n(\phi-\psi_{n}))\right]$ (1)
@@
At high transverse momentum ($p_{T}$ >> 10 GeV), hadron production in ion-ion (AA) collisions shows significant suppression compared to expectations based on proton-proton ($p$ + $p$) collisions, a phenomenon understood as jet quenching [36, 60]. In this process, high-energy partons lose energy through both radiative and collisional interactions within the QGP [60, 54, 30]. These high-$p_{T}$ hadrons and their associated jets also exhibit non-zero azimuthal anisotropy, even though they fall outside the typical regime where the hydrodynamic flow would apply [67, 2, 13]. Instead, these anisotropies are understood to arise from the spatial inhomogeneities in the QGP, with jet quenching effects being more pronounced for partons traveling through longer paths in the QGP [38, 66, 47]. The pathlength is minimized when $\phi$ and $\psi_{n}$ are approximately aligned, leading to less jet attenuation. In contrast, when $\phi$ and $\psi_{n}$ are approximately orthogonal, the pathlength is maximized, which leads to increased jet attenuation.
@@
Both low- and high-$p_{T}$ hadrons share a standard orientation of their anisotropies, correlating with the geometry of the colliding nucleons. A long-standing theoretical challenge has been to simultaneously explain both high-$p_{T}$ suppression and the corresponding azimuthal anisotropy observed in A+A collisions, and several models have been proposed to address this issue [66, 57, 58, 28, 42, 27, 74]. Previous measurements of jet differential azimuthal anisotropies $v_{n}^{\mathrm{jet}}$ have used an event plane method where the event plane is defined by soft particles  [2, 17, 52]. Given observations in the soft sector, these jet anisotropy coefficients should also be sensitive to the measurement technique. Some studies indicate that the difference could be up to 10% [24, 76, 77], which is on the order of the uncertainty on measurements by ATLAS [5].
@@
Measurements from smaller collision systems, such as $p$ + $p$ and pp+Pb at the LHC [3, 48, 49, 12, 35, 4], and pp+Au, dd+Au, and 3He+Au at RHIC [23, 10], have also revealed significant azimuthal anisotropies at low-$p_{T}$, with patterns similar to those observed in heavy-ion collisions. These findings suggest that even smaller systems may produce short-lived droplets of the QGP; indeed, hydrodynamic models have successfully described the low-$p_{T}$ behavior in such systems. However, when it comes to high-$p_{T}$ behavior in smaller collision systems, the expected signatures of jet quenching have not been observed. While some measurements indicate that no jet suppression at high transverse momentum in pp+Pb and dd+Au collisions [20, 16, 6, 14], other studies indicate there may be some suppression [9]. No alterations in the momentum balance of dijets or hadron-jet pairs have been detected.
The ATLAS experiment has also reported non-zero azimuthal anisotropies for hadrons with $p_{T}$ up to 12 GeV, suggesting an extension of anisotropy effects into the high-$p_{T}$ regime typically associated with jet quenching. It is unclear if differential jet quenching contributes to the observed high-$p_{T}$ anisotropies [6]. This leaves two related unresolved puzzles: the lack of jet quenching observed in pp+A collisions and the mechanism that produces high-$p_{T}$ hadron anisotropies in the absence of jet quenching.
@@
Further investigation into the mechanism responsible for high-$p_{T}$ hadron anisotropies requires two key approaches: (i) comparing data models with both theory and experiment using consistent techniques and (ii) expanding measurements to incorporate multi-particle correlations. Discrepancies between event plane and cumulant methods can reach up to 10% [24, 76, 77, 8]. These variations can complicate the interpretation of experimental measurements made with the event plane method. Several factors, such as long-range non-flow effects and initial/final state fluctuations, can influence the flow correlations measured with the $2$-particle correlation method, potentially amplifying the observed flow signal. Therefore, applying $2$- and multi-particle cumulant methods in data analysis and theoretical calculations is crucial for a deeper understanding of the mechanisms behind high-$p_{T}$ hadron anisotropies.
@@
In this work, we demonstrate the feasibility of measuring the jet azimuthal anisotropies $v_{n}^{\mathrm{jet}}$ with 22- and $4$-particle cumulants in Au+Au collisions simulated with Pythia-8 [68, 69] jets merged with a TennGen [44, 56, 43] background.
@@ II.1 Simulation
Events are simulated in TennGen [44, 43, 55], a data-driven background generator designed to reproduce correlations arising from flow. In this model, the $n^{\text{th}}$-order event plane is set to zero for even values of $n$, while for odd values of $n$, the event plane is chosen randomly. TennGen multiplicities and charged particle ratios match experimentally measured values [7, 11]. Similar constraints are also applied to the transverse momentum spectra of charged pions, kaons, and protons through random sampling from an initial distribution fitted to a Blast Wave [62, 65] distribution. In TennGen, the generated particle $v_{n}$($\eta$) is sampled from a flat distribution, an acceptable approximation for the region $|\eta|<1.1$. The single particle $v_{n}(p_{T})$ are modeled using a polynomial fit to the measured $v_{n}(p_{T})$ [21, 18]. TennGen events used in this study are tuned to Au+Au collision data at $\sqrt{s_{{}_{\mathrm{NN}}}}$ = 200 GeV for 20–30% central events. TennGen serves as a hydro-like background for jet interactions. The jet signal is generated using Pythia-8 [69] with the Monash Tune 13 [70] and a minimum hard scattering transverse momentum cutoff of $\hat{p}_{T}^{\text{min}}>5$ GeV.
The leading PYTHIA jets are realigned to simulate a $\frac{dN_{\text{jet}}}{d\phi_{\text{jet}}}$ distribution with non-zero differential azimuthal anisotropies $v_{n}^{\mathrm{jet}}$. Previous measurements of $v^{\text{jet}}_{n}$ by the ATLAS collaboration indicate there may be a nonzero dependence on jet $p_{T}$  [17, 2, 5]. To ensure our method is robust to such dependence, we consider several functional forms to model a wide range of momentum dependence on the jet azimuthal anisotropies $v_{n}^{\mathrm{jet}}$ $(p_{T})$. These include three $p_{T}$ independent constant values of $v^{\text{jet}}_{n}$, a $v^{\text{jet}}_{n}(p_{T})$ which increases with $p_{T}$, and a physically unrealistic sinusoidal dependence on $p_{T}$. We are confident that any expected dependence of $v^{\text{jet}}_{n}$ on jet $p_{T}$ in experimental data will be less extreme than the latter. Leading PYTHIA jets are realigned to simulate nonzero $v_{2}^{\mathrm{jet}}$, $v_{3}^{\mathrm{jet}}$, and $v_{4}^{\mathrm{jet}}$ according to one of the following descriptions:
$v^{\text{jet}}_{n}$ $=$ $\alpha_{n},$
$v^{\text{jet}}_{n}$ $=$ $2\alpha_{n},$
$v^{\text{jet}}_{n}$ $=$ $5\alpha_{n},$
$v^{\text{jet}}_{n}$ $=$ $\alpha_{n}+\frac{\beta}{2}p_{T},$
$v^{\text{jet}}_{n}$ $=$ $\alpha_{n}\left[2+\cos{\beta p_{T}}\right],$ (2)
where $\alpha_{n}$ is a nominal scale for each harmonic $n$, and $\beta$ is constant parameter for all $n$. The values used in this study are $a_{2}=4\%,a_{3}=1\%$, $a_{4}=0.5\%$, and $\beta=0.2$. Figure 1 shows the five $v^{\text{jet}}_{n}$ used in the realignment of leading PYTHIA jets. A total of $10$ million TennGen +Pythia-8 events are generated for each sample. The simulation package developed to generate these events is publicly available [JetVnCumulant].
Figure 1: The dependence of $v_{n}^{\mathrm{jet}}$ on jet $p_{T}$ for functions used to realign leading PYTHIA jets in Pythia-8 +TennGen events. These include three $p_{T}$ independent constant values of $v^{\text{jet}}_{n}$ (red, black, and orange dotted lines), a $v^{\text{jet}}_{n}(p_{T})$ which increases with $p_{T}$ (blue dotted line), and a physically unrealistic sinusoidal dependence on $p_{T}$ (green dotted line). The $y$-axis is scaled by $a_{n}$.
@@
The merged Pythia-8 jet constituents and TennGen charged particles are clustered into anti-$k_{\mathrm{t}}$ jets with jet resolution parameter $R=0.4$ using the energy recombination scheme implemented with fastjet-3 [34]. The charged final-state hadrons from Pythia-8 p+p events are clustered separately before merging with TennGen, which defines the truth jet momentum. The $p$ + $p$ event is used if there is at least one jet with $p_{\mathrm{T},\mathrm{jet}}$ >> 10 GeV. Ghost particles with $p_{T}^{\text{ghost}}=1$ MeV are used to estimate the area of clustered merged jets as described in the FastJet manual [34].
@@
The momentum of the merged jets is corrected using the jet-multiplicity background subtraction method [56]
$p_{T,jet}^{\text{corr, N}}=p_{T,jet}^{\text{raw}}-\rho_{\text{mult}}(N_{\text{tot}}-\langle N_{\text{sig}}\rangle),$ (3)
@@
where $N_{\text{tot}}$ is the observed number of particles within the jet, $\langle N_{\text{Sig}}\rangle$ is approximated by the average number of Pythia-8 particles in a TennGen +Pythia-8 jet of a given matched uncorrected jet momentum bin, and $\rho_{\text{Mult}}$ is the mean transverse momentum per background particle. Merged jet candidates satisfying $p_{T,jet}^{\text{corr}}>5$ GeV and $A_{\text{jet}}>0.6\pi R^{2}$ [72] are geometrically matched back to the jet axis of the Pythia-8 signal jet. The closest merged jet is considered a match if $\Delta R<0.75R$ and the corresponding Pythia-8 jet momentum $p_{T,jet}^{\text{PYTHIA}}$ is taken to be the truth momentum.
II.2 Cumulant Method
@@
The traditional cumulant method [29, 46], is used to construct the $v_{n}\{2k\}(p_{T})$, where $k$ is a positive integer. In this work, the $v^{\text{jet}}_{n}\{2k\}$ is given as:
[…]
$\langle\langle 2^{\prime}_{n}\rangle\rangle$ $=$ $\langle\langle e^{\textit{i}n(\psi^{A}_{1}-\varphi^{C}_{2})}\rangle\rangle,$
$\langle\langle 2_{n}\rangle\rangle$ $=$ $\langle\langle e^{\textit{i}n(\varphi^{C}_{1}-\varphi^{C}_{2})}\rangle\rangle,$ (4)
[…]
$\langle\langle 4^{\prime}_{n}\rangle\rangle$ $=$ $\langle\langle e^{\textit{i}n(\psi^{A}_{1}+\varphi^{C}_{2}-\varphi^{C}_{3}-\varphi^{C}_{4})}\rangle\rangle$
- $2~\langle\langle e^{\textit{i}n(\varphi^{C}_{1}-\varphi^{C}_{2})}\rangle\rangle\langle\langle e^{\textit{i}n(\psi^{A}_{1}(p_{T})-\varphi^{C}_{2})}\rangle\rangle,$
$\langle\langle 4_{n}\rangle\rangle$ $=$ $\langle\langle e^{\textit{i}n(\varphi^{C}_{1}+\varphi^{C}_{2}-\varphi^{C}_{3}-\varphi^{C}_{4})}\rangle\rangle$ (5)
- $2~\langle\langle e^{\textit{i}n(\varphi^{C}_{1}-\varphi^{C}_{2})}\rangle\rangle\langle\langle e^{\textit{i}n(\varphi^{C}_{1}-\varphi^{C}_{2})}\rangle\rangle,$
@@
where $\varphi^{C}$ is the azimuthal angle of particles in the region $C$, $\psi^{A}$ is the azimuthal angle of jets in the region $A$, and $\langle\langle 2_{n}\rangle\rangle$ is the event averaged $2$-particle correlations of order $n$. The prime () symbol is used to differentiate the particle of interest correlation from the reference correlation. The integrated $\varphi$ is taken from particles labeled as reference flow particles in region $C$ with $3.1<|\eta_{C}|<5.1$. The $\psi$ is the particle of interest (i.e., jet) azimuthal angle defined within region $A$ with $|\eta_{\text{A}}|<1.1-R$. More details about the traditional cumulant method can be found in Appendix A.
@@
The differential azimuthal anisotropies for jets $v^{\text{jet}}_{n}\{2k\}$ are binned as a function of uncorrected jet momentum. Jet measurements with $p_{T}$ dependencies are subject to bin-by-bin migration from smearing due to finite resolution in the uncorrected measurement. This smearing can skew the event-averaged differential particle correlations if the functional form of $v^{\text{jet}}_{n}$ is dependent on jet momentum. If the $v^{\text{jet}}_{n}$ are independent of jet momenta, then this smearing will not change the results. Measurements of $v^{\text{jet}}_{n}$ indicate there may be a nonzero dependence on $p_{\mathrm{T},\mathrm{jet}}$, meaning that measurements using a cumulant method will need to be unfolded [17, 2, 5].
We unfold the reconstructed differential single-event $n^{\mathrm{th}}$-order $2k$-particle correlations using the Bayesian unfolding method [37] in RooUnfold 2.0.0 [32]. From Equations II.2 II.2, we define the single-event $n^{\mathrm{th}}$-order $2k$-particle correlations, for use in constructing a response matrix. That is
$\langle 2_{n}\rangle$ $=$ $\langle e^{\textit{i}n(\varphi^{C}_{1}-\varphi^{C}_{2})}\rangle,$
$\langle 2^{\prime}_{n}\rangle$ $=$ $\langle e^{\textit{i}n(\psi^{A}_{1}-\varphi^{C}_{2})}\rangle,$ (6)
$\langle 4_{n}\rangle$ $=$ $\langle e^{\textit{i}n(\varphi^{C}_{1}+\varphi^{C}_{2}-\varphi^{C}_{3}-\varphi^{C}_{4})}\rangle-2~\langle 2_{n}\rangle^{2},$
$\langle 4^{\prime}_{n}\rangle$ $=$ $\langle e^{\textit{i}n(\psi^{A}_{1}+\varphi^{C}_{2}-\varphi^{C}_{3}-\varphi^{C}_{4})}\rangle-2~\langle 2_{n}\rangle\langle 2^{\prime}_{n}\rangle,$ (7)
where $\varphi^{C}$ is the azimuthal angle of particles in the region $C$, $\psi^{A}$ is the azimuthal angle of jets in the region $A$, binned in reconstructed jet $p_{T}$. Again, the prime () symbol is used to differentiate the jet azimuthal correlations from the reference particle correlations.
Pythia-8 jets (truth jets) are matched to Pythia-8 +TennGen jets (reconstructed jets) and the momentum and azimuthal angle of the Pythia-8 jet is taken as the truth momentum ($p_{T,jet}^{\text{truth}}$) and truth azimuthal angle ($\phi_{\text{jet}}^{\text{truth}}$). A reconstructed-level jet is considered a match to the closest truth-level jet in $\eta,\phi$ space if the separation is $dR<\frac{3}{4}R$. From the matched truth jet, the truth-level single event correlation $\langle(2k)^{\prime}_{n}\rangle^{\text{truth}}$ is calculated. There are no matching criteria imposed between the reconstructed-level $\langle(2k)^{\prime}_{n}\rangle^{\text{reco}}$ and the truth-level $\langle(2k)^{\prime}_{n}\rangle^{\text{truth}}$. Both single-event correlations are calculated using the same $p_{T,jet}$ binning with a bin-width of 5GeV.
@@
We construct a four-dimensional response matrix, populated for bin migration of the yield as a function of the reconstructed-level $\langle(2k)^{\prime}_{n}\rangle^{\text{reco}}\otimes p_{T,jet}^{reco}$ and the matched truth-level $\langle(2k)^{\prime}_{n}\rangle^{\text{truth}}\otimes p_{T,jet}^{truth}$. We then unfold our reconstructed jet differential single-event $n^{\mathrm{th}}$-order $2k$-particle correlations. The unfolding procedure is repeated until the change in $\chi^{2}$ between the unfolded and truth $\langle(2k)^{\prime}_{n}\rangle$ becomes less than the uncertainties of the measured $\langle(2k)^{\prime}_{n}\rangle$. This typically occurs between 3-5 iterations, and more iterations beyond this threshold would yield diminishing increases in unfolded $\langle(2k)^{\prime}_{n}\rangle$ resolution. The average value of the resulting unfolded single-event correlation is $\langle(2k)^{\prime}_{n}\rangle^{\text{corr}}$ is calculated for each jet momentum bin and used to compute $v^{\text{jet}}_{n}(p_{T})$, according to Equations $II.2$.